# Saving SWE-Bench: A Benchmark Mutation Approach for Realistic Agent Evaluation

### v1

@@ IV Experimental Setup
IV-A Benchmarks
We mutate the following two bug-fixing benchmarks:
@@ IV-A Benchmarks
Multi-SWE-Bench (TypeScript): We mutate the TypeScript subset of Multi-SWE-Bench. We chose to focus only on the TypeScript subset (224 examples) due to the size of the overall dataset. We call the mutated version of this benchmark Multi-SWE-Bench-Mutated (TypeScript).
For each of the above benchmarks, we apply the mutation prompt in Figure 12 in the Appendix and generate multiple possible mutations for each problem statement. For each example, we randomly pick one from the set of possible mutations.
IV-B Agent & Model Configurations
We use the open-source agent OpenHands [18] for our evaluation. We use OpenHands’ default prompting for bug fixing tasks, with the only exception being the change in wording from Python to C#, for our C# benchmarks. We evaluate the agent against the following three commonly used LLMs models: GPT-4.1, Claude Sonnet 3.7 and Claude Sonnet 4.
@@ V Results
V-A Impact on Success Rate
V-A1 SWE-Bench Verified Results
Table I shows the results of running OpenHands agent against baseline and mutated SWE-Bench Verified. We see that the mutation of SWE-Bench Verified problems from formal GitHub issues to realistic user queries results in a substantial performance degradation across all agent-model combinations. For all models, the agent suffers a 20-40% drop in relative success rates compared to baseline when it encounters realistic user queries instead of a longer GitHub issue style problem description.
@@ V-A1 SWE-Bench Verified Results
V-A2 SWE-Bench C# Results
In contrast to the Python benchmark results, mutation of our internal C# benchmark yields a much smaller performance impact, as seen in Table I. The smaller absolute performance drops (2-5% versus 13-16% in SWE-Bench Verified) in agent-model performances could be for a number of reasons. One possible reason would be that these LLMs are more adept at writing python than C#. This would make sense as Python repos are more abundant than C#, so there is more training data available for these models. Not just model, agents could also be overfitted to Python development, since they were built with Python benchmarks like SWE-Bench in mind. Another reason could be LLMs overfitting to popular benchmarks like the ones in SWE-Bench Verified specifically.
V-A3 Multi-SWE-Bench (TypeScript) Results
Table I (c) shows the results for running the agent on the TypeScript subset of Multi-SWE-Bench with and without mutations. Once again we see drops in agent performance for all 3 models when the bug descriptions are mutated. In fact, the relative performance degradation for TypeScript benchmark exceeds the degradation observed in the C# benchmark in all 3 models. Interestingly, we see the largest drop in the results for Sonnet 4. The model also performs worse compared to other models with the baseline benchmark as well.
V-B Impact on Steps & Token Usage
V-B1 SWE-Bench Verified Results
Looking at Table I, we see a rise in the number of steps taken by the agent. This is expected because we’ve taken away a lot of the details in the GitHub issue, so the agent has to deliberate for longer before arriving to a solution. However there is a drop in the token usage of the agent when we mutate SWE-Bench Verified. This is interesting because we would expect the agent to take more tokens proportional to the rise in the number of steps.
One more interesting insight into the numbers overall for all benchmarks is that Claude Sonnet 4 seems to be more verbose and takes more steps than the other two models, showing the model’s tendency to engage with the problem longer compared to the other models.
V-B2 SWE-Bench C# Results
We see that in this case the number of steps doesn’t rise quite as drastically as the SWE-Bench Verified results. Claude Sonnet 4 seems to suffer the highest increase in the number of steps, followed by Sonnet 3.7. Interestingly, the number of steps taken by GPT-4.1 actually go down, albeit slightly. We should note that the performance of the agent with this model is also lower, so the drop in steps may not entirely be a positive sign.
@@ V-B2 SWE-Bench C# Results
V-B3 Multi-SWE-Bench (TypeScript) Results
We see the number of steps taken by the agent to solve a given task rises for all the models when mutations are applied. Similar to the success rate, the rise in steps and token usage is also the most drastic for Sonnet 4.
Similar to SWE-Bench C#, Sonnet 4 also consumes a lot more tokens and steps for TypeScript bug-fixing tasks unlike tasks in SWE-Bench Verified. GPT-4.1 also consumes a lot more steps and tokens for TypeScript compared to the other two benchmarks.
VI Limitations
@@
Language / Task Coverage: Firstly, our work focuses primarily on benchmarks for Python and C# languages. Expanding to other programming languages will help strengthen the generalizability of our approach. Additionally, the benchmarks we examine only cover bug-fixing tasks. As demonstrated in our analysis of user queries to chat-based agents, bug fixing accounts for less than 15% of all queries. Our methodology’s efficacy for other coding tasks (feature implementation, code refactoring, software testing, etc.) remains unexplored. In future work, we plan to investigate generalizing this approach to other categories of tasks that developers commonly use coding agents for.
Agent Coverage: We evaluate only one open-source agent (OpenHands) on our mutated benchmarks. While this represents a reasonable choice given that OpenHands is among the state-of-the-art research agents [3], other agents may have different capabilities and exhibit other performance patterns when used with mutated benchmarks. Furthermore, our evaluation framework focuses on single-interaction benchmarks and does not account for the iterative, conversational nature of real chat-based agents, where users can provide clarifications and additional context through follow-up exchanges.
Template Extraction & Reliance on LLM-based Analysis: Although we believe the templates we identified are comprehensive for our dataset and usage patterns, developers may communicate differently with other coding agents whose telemetry we don’t have access to. Domain-specific communication patterns may not be represented in our telemetry sample. Additionally, our LLM-based categorization, template extraction, and application approach, despite human validation, may introduce systematic biases due to our choice of model and prompt design.
@@
VII Conclusion
In this study, we have demonstrated that existing bug-fixing benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, systematically overestimate agent capabilities by up to 20%, due to a combination of their heavy reliance on formal GitHub issue descriptions, language parity and overfitting. By analyzing real-world developer interactions with chat-based coding agents and extracting templates of how user communicate bugs to a chat, we introduced a benchmark mutation methodology that transforms problems from GitHub issue style communication to a more realistic chat-style query. Our empirical results and analysis show that this approach reveals significant performance gaps in agents and agent evaluation methodologies, highlighting the importance of realistic evaluation scenarios and the risks of benchmark overfitting. We advocate for the adoption of a mutation-based evaluation for agents and private benchmarks to better reflect actual developer usage and provide a realistic measure of agent capabilities while avoiding the pitfalls of overfitting. We believe future work should explore extending this methodology to other software engineering tasks as well as domains and further refine this technique to capture the nuances of multi-turn user-agent collaboration.

### v3

@@ 4. Experimental Setup
4.1. Benchmarks
We mutate the following three bug-fixing benchmarks:
@@ 4.1. Benchmarks
Multi-SWE-Bench (TypeScript): We mutate the TypeScript subset of Multi-SWE-Bench. We chose to focus only on the TypeScript subset (224 examples) due to the size of the overall dataset. We call the mutated version of this benchmark Multi-SWE-Bench (TypeScript)-Mutated.
For each of the above benchmarks, we apply the mutation prompt in Figure 11 and generate multiple possible mutations for each problem statement. For each example, we randomly pick one mutation from the set of possible mutations generated by the model.
4.2. Agent & Model Configurations
We use the open-source agent OpenHands (Wang et al., 2024) for our evaluation. We use OpenHands’ default prompting for bug fixing tasks, with the only exception being the change in wording from Python to C#, for our C# benchmarks. We evaluate the agent against the following three commonly used LLMs models: GPT-4.1, Claude Sonnet 3.7 and Claude Sonnet 4.
@@ 4.3. Metrics
While a higher success rate indicates better chances of task completion, it often comes with increased agent runtime and computational/monetary cost (higher token usage and step count). The user often wants to balance quality and cost, so we analyzed the effects to both classes of metrics.
@@ 5. Results
5.1. Impact on Success Rate
5.1.1. SWE-Bench Verified Results
Table 1 shows the results of running OpenHands agent against baseline and mutated SWE-Bench Verified. We see that the mutation of SWE-Bench Verified problems from formal GitHub issues to realistic user queries results in a substantial performance degradation across all agent-model combinations. For all models, the agent suffers a 20-40% drop in relative success rates compared to baseline when it encounters realistic user queries instead of a longer GitHub issue style problem description.
@@ 5.1.1. SWE-Bench Verified Results
5.1.2. SWE-Bench C# Results
In contrast to the Python benchmark results, mutation of our internal C# benchmark yields a much smaller performance impact, as seen in Table 1. The smaller absolute performance drops (2-5% versus 13-16% in SWE-Bench Verified) in agent-model performances could be for a number of reasons. One possible reason could be that LLMs or the agent are overfitting to popular benchmarks like SWE-Bench Verified. Another could be that the agent’s performance on C# isn’t very high to begin with. This could be because LLMs are simply more adept at writing python than C#. This would make sense as Python repos are more abundant than C# (Holdsworth, [n. d.]), so there is more training data available for these models. Not just model, agents could also be overfitted to Python development, since they were built with Python benchmarks like SWE-Bench in mind.
5.1.3. Multi-SWE-Bench (TypeScript) Results
Table 1 (c) shows the results for running the agent on the TypeScript subset of Multi-SWE-Bench with and without mutations. Once again we see drops in agent performance for all 3 models when the bug descriptions are mutated. In fact, the relative performance degradation for TypeScript benchmark exceeds the degradation observed in the C# benchmark in all 3 models. Interestingly, we see the largest drop in the results for Sonnet 4. The model also performs worse compared to other models with the baseline benchmark as well.
5.2. Impact on Steps & Token Usage
5.2.1. SWE-Bench Verified Results
Looking at Table 1, we see a rise in the number of steps taken by the agent. This is expected because we’ve taken away a lot of the details in the GitHub issue, so the agent has to deliberate for longer or search for relevant code before arriving to a solution. However, there is a drop in the token usage of the agent when we mutate SWE-Bench Verified. This is interesting because we would expect the agent to take more tokens proportional to the rise in the number of steps. We believe that this is because the agent has fewer details to work with — each step in the trajectory becomes less verbose.
One more interesting insight into the numbers for all benchmarks is that Claude Sonnet 4 seems to be more verbose and takes more steps than the other two models, showing the model’s tendency to engage with the problem longer compared to the other models.
5.2.2. SWE-Bench C# Results
We see that in this case, the number of steps doesn’t rise quite as drastically as the SWE-Bench Verified results. This could again be an artifact of overfitting. Claude Sonnet 4 seems to suffer the highest increase in the number of steps, followed by Sonnet 3.7. Interestingly, the number of steps taken by GPT-4.1 actually goes down, albeit slightly. We should note that the performance of the agent with this model is also lower, so the drop in steps may not entirely be a positive sign.
@@ 5.2.2. SWE-Bench C# Results
5.2.3. Multi-SWE-Bench (TypeScript) Results
We see the number of steps taken by the agent to solve a given task rises for all the models when mutations are applied. Similar to the success rate, the rise in steps and token usage is also the most drastic for Sonnet 4. Looking more closely at the agent trajectories, we notice that for both mutated and unmutated benchmarks the agent exhausts max steps in 3 times as many instances for Sonnet 4, compared to Sonnet 3.7. This shows Sonnet 4 tends to struggle with TypeScript bug fixing in general.
Similar to SWE-Bench C#, Sonnet 4 also consumes a lot more tokens and steps for TypeScript bug-fixing tasks, unlike tasks in SWE-Bench Verified. GPT-4.1 also consumes a lot more steps and tokens for TypeScript compared to the other two benchmarks.
6. Limitations
@@
Language / Task Coverage: Firstly, our work focuses primarily on benchmarks for Python, C# and TypeScript languages. Expanding to other programming languages will help strengthen the generalizability of our approach. Additionally, the benchmarks we examine only cover bug-fixing tasks. As demonstrated in our analysis of user queries to chat-based agents, bug fixing accounts for less than 15% of all queries. Our methodology’s efficacy for other coding tasks (feature implementation, code refactoring, software testing, etc.) remains unexplored. In future work, we plan to investigate generalizing this approach to other categories of tasks that developers commonly use coding agents for.
Agent Coverage: We evaluate only one open-source agent (OpenHands) on our mutated benchmarks. While this represents a reasonable choice given that OpenHands is among the state-of-the-art research agents (bench Team, 2025), other agents may have different capabilities and exhibit other performance patterns when used with mutated benchmarks. Furthermore, our evaluation framework focuses on single-interaction benchmarks and does not account for the iterative, conversational nature of real chat-based agents, where users can provide clarifications and additional context through follow-up exchanges.
Template Extraction & Reliance on LLM-based Analysis: Although we believe the templates we identified are comprehensive for our dataset and usage patterns, developers may communicate differently with other coding agents with different design and UI elements, whose telemetry we don’t have access to. Domain-specific communication patterns may not be represented in our sample. Additionally, our LLM-based categorization, template extraction, and application approach, despite human validation, may introduce systematic biases due to our choice of model and prompt design.
@@
7. Conclusion
In this study, we have demonstrated that existing bug-fixing benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, Multi-SWE-Bench, systematically overestimate agent capabilities by 50% over baselines, due to a combination of their heavy reliance on formal GitHub issue descriptions, language parity and overfitting. By analyzing real-world developer interactions with chat-based coding agents and extracting templates of how user communicate bugs to a chat, we introduced a benchmark mutation methodology that transforms problems from GitHub issue style communication to a more realistic chat-style query. Our empirical results and analysis show that this approach reveals significant performance gaps in agents and agent evaluation methodologies, highlighting the importance of realistic evaluation scenarios and the risks of benchmark overfitting. We advocate for the adoption of a mutation-based evaluation for agents and private benchmarks to better reflect actual developer usage and provide a realistic measure of agent capabilities while avoiding the pitfalls of overfitting. We believe future work should explore extending this methodology to other software engineering tasks as well as domains and further refine this technique to capture the nuances of multi-turn user-agent collaboration.